An aircraft on its take off run has a steady acceleration of 3m/s^2.
What velocity does the aircraft gain in 4s?
If the aircraft passes one post on the runway at a velocity or 20m/s, what is it’s velocity 8s later?
1
Expert's answer
2019-02-14T11:31:07-0500

The velocity change

Δv=aΔt=3 m/s^2 ×4 s=12 m/s

The final velocity

v_f=v_i+at=20 m/s+2 m/s^2 ×8 s=36 m/s